OpenStudy (anonymous):

What is the domain and range of the quadratic equation y = -x2 - 14x - 52? I know the domain is all real numbers, but I'm stumped on how to find the range. Help?

OpenStudy (anonymous):

You could graph it.

OpenStudy (anonymous):

I can't -- I don't have any tools to graph with. How do I find it without graphing is what I need to know. :[

OpenStudy (anonymous):

domain is all real numbers because it is a polynomial

OpenStudy (anonymous):

any polynomial has domain all real numbers, you never have to worry

OpenStudy (anonymous):

I know the domain is all real numbers :P I just need to know how I can find the range without graphing, haha.

OpenStudy (anonymous):

as for the range, this is a parabola that opens down

OpenStudy (anonymous):

Find the vertex.

OpenStudy (anonymous):

to the range will be from minus infinity to the highest point on the parabola, which is easy to find

OpenStudy (anonymous):

Satellite knows what's up.

OpenStudy (anonymous):

the polynomial is \[y = -x2 - 14x - 52\] and to find the highest point, locate the second coordinate of the vertex. use \[x=-\frac{b}{2a}\]

OpenStudy (anonymous):

in this case \[a=-1,b=-14,-\frac{b}{2a}=-\frac{-14}{-2}=-7\]

OpenStudy (anonymous):

that is the first coordinate, you need the second coordinate which you find by replacing x in the expression \[y = -x^2 - 14x - 52\] by -7

OpenStudy (anonymous):

you get \[y=-49+98-52=-3\]

OpenStudy (anonymous):

so the vertex is (-7,-3), the maximum value of the equation is -3 and the range is \[(-\infty,-3]\]

OpenStudy (anonymous):

let me know if steps are clear

OpenStudy (anonymous):

they are :D Thanks so much!

OpenStudy (anonymous):

yw don't forget \[-\frac{b}{2a}\] is the magic formula